### v4

# Dynamic EventNeRF: Reconstructing General Dynamic Scenes from Multi-view RGB and Event Streams

Abstract
Volumetric reconstruction of dynamic scenes is an important problem in computer vision. It is especially challenging in poor lighting and with fast motion. This is partly due to the limitations of RGB cameras: To capture frames under low lighting, the exposure time needs to be increased, which leads to more motion blur. In contrast, event cameras, which record changes in pixel brightness asynchronously, are much less dependent on lighting, making them more suitable for recording fast motion. We hence propose the first method to spatiotemporally reconstruct a scene from sparse multi-view event streams and sparse RGB frames. We train a sequence of cross-faded time-conditioned NeRF models, one per short recording segment. The individual segments are supervised with a set of event- and RGB-based losses and sparse-view regularisation. We assemble a real-world multi-view camera rig with six static event cameras around the object and record a benchmark multi-view event stream dataset of challenging motions. Our work outperforms RGB-based baselines, producing state-of-the-art results, and opens up the topic of multi-view event-based reconstruction as a new path for fast scene capture beyond RGB cameras.
1 Introduction
Spatiotemporal (or 4D) reconstruction of a non-rigid scene allows re-rendering it from novel viewpoints. This is a long-standing and challenging problem in computer vision [63, 57]. Until recently, the great majority of methods used RGB frames as input data. However, RGB frames contain motion blur for fast objects and become noisier the less light is available in the scene. This has motivated the exploration of event cameras [64, 5, 10] as a data source for this task, which have greater dynamic range and superior temporal resolution. Since event cameras are by far not as commonplace as RGB cameras and thus still more expensive, previous event-based non-rigid spatiotemporal capture techniques [25, 48, 65, 60, 59, 28] are mostly monocular. Using only one camera can be expected to limit the ability to capture general dynamic scenes, especially when handling fast motion, large deformations and occlusions [11].
In this work, we explore the use of multi-view event data for event-based 4D reconstruction of non-rigid motion. Since a multi-view event camera setup is, of course, more complex than a monocular setup, we believe that investigating this uncharted territory can prove a helpful guideline in deciding whether the added cost is worth the gains in quality.
While previous work [49, 13, 19] addressed the event-based reconstruction of static scenes, this work reconstructs dynamic scenes, i.e. yields different 3D models for different time points. A common approach in reconstructing dynamic scenes is to explain each state as a deformation of a canonical volume [39, 40, 52]. However, this tends to limit both the range and kinds of reconstructed motion [53, 11]. To address these challenges, we train a sequence of cross-faded time-conditioned NeRF models, each representing a short recording segment. This way, smaller motions can be learned locally with consistent quality, even if the full recording is long. Moreover, having no explicit deformation models, we can reconstruct general motion.
To evaluate our method, we both render synthetic data and obtain real data by setting up a multi-view rig of six event cameras, with which we record a benchmark dataset of various scenes, subjects and motions under lighting conditions ranging from “dimly lit” to “very dark” (LABEL:fig:teaser). We compare our method to RGB-based baselines trained on blurry RGB videos and RGB videos reconstructed from events. Our method significantly outperforms all baselines and produces state-of-the-art results, while providing a continuous reconstruction of the scene. It demonstrates that operating directly on events instead of reconstructed RGB images in challenging conditions results in higher novel-view rendering accuracy. To summarise, our contributions are as follows:
  • 1)
    We present Dynamic EventNeRF, the first method for learning general 4D scenes from sparse multi-view event streams and sparse blurry RGB frames;
  • 2)
    We demonstrate the usage of time-conditioned NeRF models, our blended multi-segment approach, an event-based supervision scheme, and fast event accumulation with decay to learn the sequences;
  • 3)
    We provide multi-view static camera real and synthetic datasets for 4D reconstruction from event streams.
Our datasets and the source code are available online [^1] https://4dqv.mpi-inf.mpg.de/DynEventNeRF/.
2 Related Works
NeRF for non-rigid scenes with RGB inputs and sparse views
Novel view synthesis has seen recent progress [26, 51, 23, 27, 17], with methods based on Neural Radiance Fields (NeRF) [27] being especially successful. NeRF has been adapted for dynamic scene reconstruction, e.g. head avatars [45], full-body avatars [41, 22, 12], hands [32] and other domains. Other techniques target general scene reconstruction, which can be divided into monocular [43, 40, 52, 20] and multi-view methods [9, 4, 50, 42, 58, 54]. While the monocular methods can reconstruct the input video with high fidelity, the ambiguities of monocular input make it hard for them to learn the correct geometry, greatly limiting how far novel views can deviate from the training viewpoints. Moreover, many monocular methods greatly depend on the camera moving faster than the object, constraining the scenes they can reconstruct [11]. As we capture fast motion, moving the camera even faster is not a viable option. Our method uses a fixed camera multi-view setup. Multi-view methods generalise better to novel views, but are still limited by their RGB input, which is prone to motion blur and noise, especially with fast motion or low lighting conditions. Instead, we use event cameras, possessing both high temporal resolution and high dynamic range. To train NeRF models, one typically needs dozens of ground-truth RGB views, but in our setting, we only have six views available. Follow-up works of NeRF aim at enabling few-view reconstruction, using extensive regularisation [36], pre-training [62, 33], semantic consistency [15], or depth priors [34, 47]. Our approach (Sec. 4.3) uses a combination of regularisers and volume clipping, without a need for pretraining or additional priors, and is applicable to general scenes.
Event-based vision for dynamic scenes and 3D reconstruction
Event cameras have been used to reconstruct non-blurry RGB videos of fast motion: Some methods [46, 38] use a learning-based approach to convert events into RGB frames, while others [37, 55, 56] combine events and input RGB frames and use the former to deblur and interpolate the latter. However, all of them only support 2D reconstruction. Methods based on 3D meshes and parametric models allow tracking of full bodies [59, 65, 2], hands [48, 35, 60, 16, 29] and general templates [35, 60], but do not reconstruct appearance and are not applicable to general scenes, which is what our Dynamic EventNeRF is aiming for.
For general static scenes, the state of the art is represented by a new class of NeRF-based methods that allow dense 3D reconstruction and novel-view synthesis using event streams: EventNeRF [49] and Ev-NeRF [13] supervise NeRF using accumulated windows of events, E-NeRF [19] supervises each event directly, Robust E-NeRF[24] extends it to account for the pixel refraction and noise, E2NeRF [44] and Ev-DeblurNeRF [3] use events to enhance and deblur RGB-based scene reconstructions. These methods model static scenes using data captured with a single moving camera, while our Dynamic EventNeRF reconstructs general dynamic scenes using a multi-view fixed camera setup.
DE-NeRF [25] reconstructs dynamic scenes with monocular event streams and RGB frames from a moving camera through deformations of a canonical volume. As discussed earlier, such a setting and approach significantly limit the kind of scenes and motions that can be reconstructed. In contrast, our Dynamic EventNeRF uses a multi-view fixed camera setup and is not based on canonical volumes, allowing it to reconstruct general dynamic scenes.
EvDNeRF [1] allows to synthesise events from novel viewpoints using multi-view event streams through the deformation of a canonical volume. It does not model the appearance, needs as many as 18 views, and the deformation-based approach significantly limits the scenes and motions that can be reconstructed. In contrast, our Dynamic EventNeRF allows to reconstruct general dynamic scenes directly in RGB space while using as few as three views.
3 Background
3.1 Event Camera Model and Deblurring
Instead of capturing dense RGB frames at regular intervals, event cameras operate asynchronously per individual pixel: When a certain pixel gets brighter or darker by a certain threshold, it emits an event, which is a tuple $(t,x,y,p)$ of its timestamp $t$, pixel position $x,y$, and polarity $p\in\{-1,+1\}$. The event can be interpreted as the assertion
$\log I_{x,y}(t)-\log I_{x,y}(t^{\prime})=pC_{p},$ (1)
where $\log I_{x,y}(\tau)$ is the intensity of pixel $(x,y)$ at time $\tau$ and $t^{\prime}<t$ is the timestamp of the previous event in that pixel. We call $C_{\mathrm{+1}},C_{\mathrm{-1}}>0$ the event generation thresholds. Event-based Single Integral (ESI) [49] connects the difference between logarithmic intensities on its left side and the accumulated events on the right side by summing Eq. 1 over time. ESI can be further extended to Event-based Double Integral (EDI) [37] by averaging the intensity over the exposure period, to model motion blur caused by the traditional camera shutter. Given events $\mathbf{E}$ and the blurry image $\mathbf{B}$, one can solve a system of equations generated by EDI for the unknown instantaneous intensity image $\mathbf{I}$, thus deblurring $\mathbf{B}$. This way, EDI can be used for deblurring the RGB stream through events. To synthesise arbitrary instantaneous frames $\mathbf{I}(t)$, one could accumulate events on top of the closest recovered deblurred frame $\mathbf{I}(t_{0})$ with ESI. To improve the speed of deblurring and frame synthesis, Lin et al. propose Fast EDI [21], which uses the same event camera model and a special data structure to accelerate the queries.
3.2 EventNeRF
Rudnev et al. [49] represent a static scene as an MLP predicting point density $\sigma(\mathbf{p})\in\mathbb{R}$ and colour $\mathbf{C}(\mathbf{p},\mathbf{\theta})$ as functions of its position $\mathbf{p}\in\mathbb{R}^{3}$ and view directions $\theta\in\mathbb{R}^{2}$. Novel views are obtained through volumetric rendering. First, they associate a ray $\mathbf{r}^{(x,y,t)}(d)\in\mathbb{R}^{3}$ with every pixel $(x,y)$ that they render from the viewpoint at time $t$. Then they randomly sample them at depths $\{d_{i}\}_{i=1}^{N_{\textrm{samples}}}$: $\{\mathbf{r}^{(x,y,t)}_{i}\}_{i=1}^{N_{\textrm{samples}}}$. And finally, they integrate them into the pixel colour as:
$\mathbf{\hat{C}}_{x,y,t}=\sum_{i=1}^{N_{\textrm{samples}}}(1-\exp(-\sigma_{i}))T_{i}\mathbf{C}_{i},$ (2)
where $T_{i}=\exp\left(-\sum_{j=1}^{N_{\textrm{depth}}-1}\sigma_{j}(d_{j+1}-d_{j})\right)$ is the accumulated transmittance, $d_{j}$ is the distance of the $j^{\textrm{th}}$ ray sample from the ray origin, $\sigma_{i}=\sigma\left(\mathbf{r}_{i}^{(x,y,t)}\right)$ and $\mathbf{C}_{i}=\mathbf{C}\left(\mathbf{r}^{(x,y,t)}_{i},\mathbf{\theta}^{(x,y,t)}\right)$ are the sample density and colour.
To supervise this model with ground-truth events, the authors use ESI (Sec. 3.1): They substitute intensity values $I_{x,y}(t)$ with the rendered predictions $\hat{\mathbf{C}}_{x,y,t}$ and apply the MSE loss between the intensity and event sides of ESI:
$\mathcal{L}_{x,y}(t_{0},t)={\|\hat{E}_{x,y}(t_{0},t)-E_{x,y}(t_{0},t)\|^{2}},$ (3)
where $\hat{E}_{x,y}(t_{0},t)=\mathbf{F}(\log\hat{\mathbf{C}}_{x,y,t}-\log\hat{\mathbf{C}}_{x,y,t_{0}})$ is the predicted difference, $E(t_{0},t)=\sum_{i:t_{0}<t_{i}\leq t}{p_{i}C_{p_{i}}\delta_{x_{i},y_{i}}^{x,y}}$ is the accumulated difference, $\mathbf{F}$ is the colour filter, $\delta_{\mathbf{a}}^{\mathbf{\hat{a}}}=1$ iff. $\mathbf{a}=\mathbf{\hat{a}}$, 00 otherwise. The performance is also much improved with Positional Encoding [27].
4 Our Dynamic EventNeRF Approach
Figure 1: Overview of Dynamic EventNeRF. We split the entire sequence into short overlapping segments (e.g., $\mathrm{seg}~{}i-1$, $\mathrm{seg}~{}i$, $\mathrm{seg}~{}i+1$, $\mathrm{seg}~{}i+2$ on the bottom-left of the figure). For each segment, we learn a time-conditioned MLP-based NeRF model. To supervise it, we first sample a random window $[t_{0},t_{1}]$ within the segment and apply a combination of the following losses: 1) Event loss, supervising predicted view differences; 2) Accumulation loss, supervising differences between a reference RGB frame and one of the predicted views; 3) RGB loss, supervising the model with the reference RGB frame, and 4) Sparsity loss, that minimises the number of opaque pixels in each predicted views. For increased stability and fast computation of these losses, we propose Fast Event Accumulation with Decay; see Sec. 4.4.
[…]
4.1 Model for One Segment
We reconstruct a 4D model of the $i^{\mathrm{th}}$ temporal interval $[t_{\mathrm{start}_{i}},t_{\mathrm{end}_{i}}]$, with $0\leq t_{\mathrm{start}_{i}}\leq t_{\mathrm{end}_{i}}\leq T_{\mathrm{end}}$, using sparse multi-view events $\{\mathbf{E}^{k}\}_{k=1}^{K}$ and the supporting instantaneous (non-blurry) RGB views $\{\mathbf{C}^{k}(t_{\mathrm{ref}_{i}}),\mathbf{C}^{k}(t_{\mathrm{ref}_{i+1}})\}_{k=1}^{K}$ at the start and the end of the interval (Fig. 1, left). For convenience of notation, we linearly map the temporal extent of the interval to $-1\leq t\leq 1$. To represent the scene, we use an MLP mapping space-time coordinates $\mathbf{p}=[x,y,z,t]^{T}$ to density $\sigma\in\mathbb{R}$ and colour $\mathbf{C}\in\mathbb{R}^{3}$, where $-1\leq x,y,z,t\leq 1$. We obtain the foreground ray colours $\mathbf{\hat{C}}^{k,\textrm{fg}}$ of the training view $k$ through volumetric rendering of $\sigma$ and $\mathbf{C}$, similar to NeRF. To obtain the final ray colours $\mathbf{\hat{C}}$, we alpha-blend the foreground and background:
$\mathbf{\hat{C}}^{k}_{x,y}=\mathbf{\hat{C}}_{x,y}^{k,\textrm{fg}}+T_{N_{\mathrm{samples}}+1}\cdot\mathbf{A}^{k}_{x,y},$ (4)
where $T_{N_{\mathrm{samples}}+1}$ represents background opacity (Eq. 2). Note that instead of learning the background with a separate model, we use the ground-truth background $\mathbf{A}^{k}$ for predictions of the view $k$. This eliminates possible artefacts and ambiguities due to imprecise background modelling.To facilitate the training, we use separate Positional Encodings for embedding time and spatial coordinates $\mathbf{p}$ (Sec. 3.2). The maximum frequency of the encoding determines how coherent the model is across the corresponding dimension. If it is zero, the model would have no variation over that dimension, and if it is too high, then the model would behave randomly due to aliasing. We choose the optimal values empirically to be 7 temporal and 14 spatial frequencies.
Using a single MLP network to represent longer periods of time would pose a number of challenges: The network would have to be sufficiently large, significantly slowing down both inference and training. Moreover, supervising the MLP at one moment of time $t$, could, as an unwanted side effect, cause artefacts at $t^{\prime}\neq t$, as both times are represented with a single network. In the next section, we overcome these issues by splitting the full sequence into smaller segments and training one model per segment.
4.2 Multi-Segment Model
Using a single MLP to represent longer sequences is challenging due to the capacity-performance trade-off and the ghosting artefacts caused by using a shared network. Therefore, we propose to split the full sequence into short segments and train one model per such sequence. To eliminate flickering caused by changing from one model to the next, we overlap the segments by $10\%$ at the starts and at the ends. In the overlapped parts, we render both models as $\mathbf{\hat{C}}_{i}(t)$ and $\mathbf{\hat{C}}_{i+1}(t)$ and cross-fade the predictions:
$\mathbf{\hat{C}}(t)=(1-\alpha(t))\mathbf{\hat{C}}_{i}(t)+\alpha(t)\mathbf{\hat{C}}_{i+1}(t),$ (5)
where $\alpha(t)=\frac{t-t_{\textrm{ov. start}}}{t_{\textrm{ov. end}}-t_{\textrm{ov. start}}}$, with $t_{\textrm{ov. start}}$ and $t_{\textrm{ov. end}}$ being the overlapping start and end times, respectively. In the next section, we describe the training procedure and all the regularisation techniques we use to adapt the method to event supervision and the sparse-view setting.
4.3 Supervision for One Segment
In this section, we describe how we supervise the model for a single segment $i$ from $t_{\mathrm{start}_{i}}$ to $t_{\mathrm{end}_{i}}$ (Fig. 1). We use sparse multi-view events $\{\mathbf{E}^{k}_{i}\}_{k=1}^{K}$ and the supporting instantaneous RGB views $\{\mathbf{C}^{k}(t_{\mathrm{ref}_{i}}),\mathbf{C}^{k}(t_{\mathrm{ref}_{i+1}})\}_{k=1}^{K}$ at the start and the end of the segment.
Event supervision
For event input, we start by randomly choosing a temporal window $[t_{0},t_{1}]$: $t_{1}$ is sampled from the uniform distribution $U[t_{\mathrm{start}_{i}},t_{\mathrm{end}_{i}}]$; $t_{0}$ is chosen such that the window duration $t_{1}-t_{0}$ is distributed uniformly from $10\%$ to $30\%$ of the segment’s length. For each temporal window, per each view $1\leq k\leq K$, we accumulate the events inside the window into $E^{k}(t_{0},t_{1})\in\mathbb{R}^{W\times H}$ according to the model described in Sec. 4.4
Similarly to EventNeRF [49], we form training batches by combining $10\%$ positive and $90\%$ negative rays jointly for each view $k$: Positive rays are randomly chosen pixels $x,y$ for which $E_{x,y}(t_{0},t_{1})\neq 0$, i.e. from regions where events occurred. Negative rays are the pixels uniformly sampled from all pixels in the view. Including them allows us to reduce noise and reconstruct uniformly coloured details. For the colour values obtained for the sampled rays, we compute an MSE loss $L_{\mathrm{event}}$ in the event camera log-space:
$\small\begin{split}&K\cdot\mathcal{L}_{\mathrm{event}}=\\ &\sum_{k=1}^{K}\mathrm{MSE}\left(E^{k}(t_{0},t_{1}),F\left(\log{\mathbf{\hat{C}}}^{k}(t_{1})-\log{\mathbf{\hat{C}}}^{k}(t_{0})\right)\right),\end{split}$ (6)
where $\mathbf{\hat{C}}^{k}(t)\in\mathbb{R}^{3}$ is the model prediction at time $t$ from view $k$, and $F(\cdot):\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{H\times W}$ is applying the Bayer filter to the RGB image, resulting in a greyscale intensity image matching the event camera colour filter.
RGB supervision
To make the model match RGB ground-truth values $\{\mathbf{C}^{k}(t_{\mathrm{ref}})\}_{k=1}^{K}$, $t_{\mathrm{ref}}\in\{t_{\mathrm{ref}_{i}},t_{\mathrm{ref}_{i+1}}\}$, we compute another MSE loss:
$\mathcal{L}_{\mathrm{RGB}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{MSE}\left(\mathbf{C}^{k}(t_{\mathrm{ref}}),{\mathbf{\hat{C}}}^{k}(t_{\mathrm{ref}})\right),$ (7)
where ${\mathbf{\hat{C}}}^{k}(t_{\mathrm{ref}})$ is the model prediction.
Accumulation loss
Eq. 7 only supervises the model at the reference RGB frame timestamps. Because of that, the information from the reference frame is not propagated well to the rest of the sequence. To mitigate this issue, we also use the accumulation loss, which connects the closest reference RGB frame $\mathbf{C}^{k}(t_{\mathrm{ref}})$, the prediction at the end of the window $\mathbf{\hat{C}}^{k}(t_{1})$, and the accumulated events $E^{k}(t_{\mathrm{ref}},t_{1})$ in between:
$\small\begin{split}&K\cdot\mathcal{L}_{\mathrm{acc}}=\\ &\sum_{k=1}^{K}\mathrm{MSE}\left(E^{k}(t_{\mathrm{ref}},t_{1}),F\left(\log{\mathbf{\hat{C}}}^{k}(t_{1})-\log{\mathbf{C}}^{k}(t_{\mathrm{ref}})\right)\right).\end{split}$ (8)
Sparse-view regularisation
For sparse-view reconstruction, we use the following combination of regularisation techniques: First, we want to prevent the model from learning artefacts in the regions where only one view observes them. These areas are near and far away from the camera, but still in its frustum. Suppose the views are distributed 360°360°360  ° around the object at roughly similar height: The intersection of all frustums can be approximated by a cylinder of radius $r$ that extends vertically from $y_{\mathrm{min}}$ to $y_{\mathrm{max}}$. We can then clip the reconstruction volume by setting $\sigma(x,y,z):=0$ wherever $x^{2}+z^{2}>r$, or $y>y_{\textrm{max}}$ or $y<y_{\textrm{min}}$. We also modify the ray depth sampling so that the samples are only chosen inside the cylinder. Following EventNeRF, we randomly offset the ray direction within a pixel to improve generalisation to novel views. This cylinder clipping and the improved sampling procedure result in a strong prior significantly reducing the ambiguity of the sparse-view reconstruction.
As stated earlier, we want our model only to reconstruct the foreground. Since we use the cylindrical bounding volume, the model does not have a way to represent the background (that is out of bounds) correctly. However, the model can still create background-coloured artefacts within the bounds, which it can also use to hide view-inconsistent artefacts. To resolve these issues, we minimise the number of rays hitting the foreground (or, equivalently, maximise the number of rays hitting the actual background) by applying a per-ray sparsity loss:
$\mathcal{L}_{\mathrm{sparsity}}(n)=\gamma_{n}\cdot\frac{1}{WH}\sum_{{x,y}=1}^{WH}\left(1-T_{N_{\mathrm{samples}}+1}^{(x,y)}\right),$ (9)
[…]
Learning high-frequency details from the start is not optimal, especially in the sparse-view setting. Thus, we implement coarse-to-fine training by applying positional encoding annealing [39] both across spatial frequencies and time. Annealing improves temporal and spatial smoothness as the method starts with a coarse representation, slowly progressing towards a finer representation.
Full loss
We combine all aforementioned losses:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\mathrm{event}}+\lambda_{2}\mathcal{L}_{\mathrm{acc}}+\lambda_{3}\mathcal{L}_{\mathrm{RGB}}+\lambda_{4}\mathcal{L}_{\mathrm{sparsity}},$ (10)
[…]
4.4 Event Accumulation
Following Sec. 3.1, we define event accumulation:
$E_{x,y}(t_{0},t_{1})=\sum_{i=i_{\mathrm{start}}}^{i_{\mathrm{end}}}\delta_{x_{i},y_{i}}^{x,y}p_{i}C_{p_{i}}b^{i_{\mathrm{end}}-i_{\mathrm{start}}-i},~{}b=0.93,$ (11)
where $\delta_{a,b}^{a^{\prime},b^{\prime}}:=1$ for $(a,b)=(a^{\prime},b^{\prime})$ and 00 otherwise; $C_{p}$ is the event generation threshold for polarity $p\in\{-1,+1\}$ and $i_{\mathrm{start}}$, $i_{\mathrm{end}}$ indicate the indices of the first and last events in $(t_{0},t_{1}]$, respectively. The term $b^{i_{\mathrm{end}}-i_{\mathrm{start}}-i}$, with $b:=0.93$ makes events that are far in the past weigh less than events that are rather recent. As we show in Appx. I of our appendix, this is necessary to make our approach robust against noise events.
In Eq. 6, we defined $\mathcal{L}_{\mathrm{event}}$, which uses $E(t_{0},t_{1})$ for random $t_{0}$ and $t_{1}$ each training iteration. As event streams could contain millions of events per second, implementing the accumulation procedure directly by processing each event on every query would be too slow to run every training iteration. To overcome this issue, we use an approach inspired by Lin et al. [21] to accelerate the queries: We treat all pixels independently and store all intermediate accumulation results in arrays $A^{x_{i},y_{i}}_{j}=E(0,t_{i})_{x_{i},y_{i}}$, and $T^{x_{i},y_{i}}_{j}=t_{i}$, where $j$ is the per-pixel event counter. Then, once we need to compute $E(t_{0},t_{1})$, we split it into two queries:
$E(t_{0},t_{1})=E(0,t_{1})-E(0,t_{0}).$ (12)
[…]
Ground truth Our method E+Dyn-NeRF B+Dyn-NeRF
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Ground truth E+FreeNeRF [61] D+Dyn-NeRF D+NR-NeRF[52]
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: For two real scenes, we compare novel views by different methods (top: “Towel T.”, bottom: “Dancing”). “E+”, “B+”, and “D+” are trained using events processed with E2VID [46], using Blurry RGB frames, and using EDI-deblurred frames [37], respectively. Dyn-NeRF represents our method supervised only with RGB. E2VID-based baselines fail due to view inconsistencies of E2VID data and its artefacts. Blurry RGB+Dyn-NeRF fails due to the blurriness and sparsity of its inputs. NR-NeRF fails due to the fast and large motion in the scenes.
Ground truth Our method E+Dyn-NeRF
Refer to caption Refer to caption Refer to caption
B+Dyn-NeRF E+FreeNeRF [61] B+FreeNeRF [61]
Refer to caption Refer to caption Refer to caption
R+Dyn-NeRF R+NR-NeRF [52]
Refer to caption Refer to caption
Figure 3: Qualitative comparisons on synthetic “Blender” sequence. ”E+”, ”B+”, and ”R+” baselines are trained with E2VID-processed events, blurry RGB frames, and original RGB frames as inputs, respectively. Dyn-NeRF is our method without event supervision using only RGB data.
5 Experiments
We compare our method to RGB-based baselines that do not use events and are trained either on blurry RGB recordings or RGB videos generated from events using E2VID [46]. We evaluate on both synthetic and real sequences. We evaluate the predicted novel views over the reconstructed sequence by computing PSNR, SSIM, and LPIPS. For the synthetic data, we use three hold-out views to obtain quantitative results.
5.1 Synthetic Dataset
[…]
5.2 Multi-view Event Camera Setup and Data
[…]
Reference RGB Full Model TensoRF-CP [7] NGP [31] HexPlane [4]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Enhanced RGB w/o multi-seg. w/o $L_{\mathrm{sparsity}}$ w/o clipping w/o decay
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
w/o $L_{\mathrm{event}}$ w/o $L_{\mathrm{acc}}$ w/o $L_{\mathrm{RGB}}$ only $L_{\mathrm{event}}$ only $L_{\mathrm{acc}}$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Ablation study on the “Sword” real sequence. Our full method produces best quality predictions. Alternative backbones [7, 31, 4] perform worse, as their grid-based structure does not allow for smooth propagation of details across time. Disabling various parts of the method reduces the prediction quality. Especially disabling event decay makes the contours of the subject more blurry. The recordings were done in the dark, as apparent from a reference RGB video that was captured with a regular camera (not used in training). As it is pitch black, we also show the enhanced version.
Ground truth TensoRF-CP [7] NGP [31] HexPlane [4]
Refer to caption Refer to caption Refer to caption Refer to caption
Full Model w/o $L_{\mathrm{event}}$ w/o $L_{\mathrm{acc}}$ w/o $L_{\mathrm{RGB}}$
Refer to caption Refer to caption Refer to caption Refer to caption
only $L_{\mathrm{event}}$ only $L_{\mathrm{acc}}$ w/o $L_{\mathrm{sparsity}}$ w/o clipping
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Ablation studies on the “Static Lego” synthetic sequence (zoomed in). NGP fails to produce correct geometry in our sparse-view setting. TensoRF-CP and HexPlane fail due to their grid-based nature preventing propagation of details through time. Disabling parts of the method results in blurrier predictions and artefacts.
Blender Dress Spheres Lego Static Lego Average
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Our Dyn-EventNeRF 27.607 806 5427.6078065427.607\,806\,5427.607 806 54 0.929 897 798 80.92989779880.929\,897\,798\,80.929 897 798 8 0.112 565 533 80.11256553380.112\,565\,533\,80.112 565 533 8 31.146 451 1331.1464511331.146\,451\,1331.146 451 13 0.950 413 2430.9504132430.950\,413\,2430.950 413 243 0.076 753 767 490.076753767490.076\,753\,767\,490.076 753 767 49 29.841 485 1929.8414851929.841\,485\,1929.841 485 19 0.924 164 174 70.92416417470.924\,164\,174\,70.924 164 174 7 0.082 664 634 440.082664634440.082\,664\,634\,440.082 664 634 44 22.415 136 3922.4151363922.415\,136\,3922.415 136 39 0.799 447 9640.7994479640.799\,447\,9640.799 447 964 0.285 815 632 60.28581563260.285\,815\,632\,60.285 815 632 6 23.943 150 4423.9431504423.943\,150\,4423.943 150 44 0.844 104 8260.8441048260.844\,104\,8260.844 104 826 0.196 650 731 10.19665073110.196\,650\,731\,10.196 650 731 1 26.990 805 9426.9908059426.990\,805\,9426.990 805 94 0.889 605 601 30.88960560130.889\,605\,601\,30.889 605 601 3 0.150 890 059 90.15089005990.150\,890\,059\,90.150 890 059 9
E2VID [46] + Dyn-NeRF 12.473 289 112.473289112.473\,289\,112.473 289 1 0.675 842 6450.6758426450.675\,842\,6450.675 842 645 0.536 400 704 10.53640070410.536\,400\,704\,10.536 400 704 1 21.900 224 5821.9002245821.900\,224\,5821.900 224 58 0.879 776 093 30.87977609330.879\,776\,093\,30.879 776 093 3 0.332 506 959 10.33250695910.332\,506\,959\,10.332 506 959 1 19.106 602 8819.1066028819.106\,602\,8819.106 602 88 0.870 612 836 90.87061283690.870\,612\,836\,90.870 612 836 9 0.271 807 938 20.27180793820.271\,807\,938\,20.271 807 938 2 18.200 985 4918.2009854918.200\,985\,4918.200 985 49 0.724 539 4470.7245394470.724\,539\,4470.724 539 447 0.451 910 169 90.45191016990.451\,910\,169\,90.451 910 169 9 15.772 116 5515.7721165515.772\,116\,5515.772 116 55 0.697 155 670 30.69715567030.697\,155\,670\,30.697 155 670 3 0.597 945 600 70.59794560070.597\,945\,600\,70.597 945 600 7 17.490 643 7217.4906437217.490\,643\,7217.490 643 72 0.769 585 338 50.76958533850.769\,585\,338\,50.769 585 338 5 0.438 114 274 40.43811427440.438\,114\,274\,40.438 114 274 4
Blurry RGB + Dyn-NeRF 23.148 831 8523.1488318523.148\,831\,8523.148 831 85 0.889 072 509 30.88907250930.889\,072\,509\,30.889 072 509 3 0.151 478 172 30.15147817230.151\,478\,172\,30.151 478 172 3 27.412 894 9927.4128949927.412\,894\,9927.412 894 99 0.915 951 936 10.91595193610.915\,951\,936\,10.915 951 936 1 0.137 360 298 90.13736029890.137\,360\,298\,90.137 360 298 9 20.538 948 0120.5389480120.538\,948\,0120.538 948 01 0.877 933 112 20.87793311220.877\,933\,112\,20.877 933 112 2 0.211 472 264 20.21147226420.211\,472\,264\,20.211 472 264 2 20.285 844 820.285844820.285\,844\,820.285 844 8 0.749 347 084 80.74934708480.749\,347\,084\,80.749 347 084 8 0.403 806 152 90.40380615290.403\,806\,152\,90.403 806 152 9 22.999 806 1922.9998061922.999\,806\,1922.999 806 19 0.819 269 518 10.81926951810.819\,269\,518\,10.819 269 518 1 0.207 238 751 90.20723875190.207\,238\,751\,90.207 238 751 9 22.877 265 1722.8772651722.877\,265\,1722.877 265 17 0.850 314 832 10.85031483210.850\,314\,832\,10.850 314 832 1 0.222 271 128 10.22227112810.222\,271\,128\,10.222 271 128 1
E2VID [46] + FreeNeRF [61] 9.069 228 1889.0692281889.069\,228\,1889.069 228 188 0.626 939 392 20.62693939220.626\,939\,392\,20.626 939 392 2 0.499 005 169 70.49900516970.499\,005\,169\,70.499 005 169 7 19.208 220 4319.2082204319.208\,220\,4319.208 220 43 0.892 123 077 70.89212307770.892\,123\,077\,70.892 123 077 7 0.286 058 750 70.28605875070.286\,058\,750\,70.286 058 750 7 17.560 102 0117.5601020117.560\,102\,0117.560 102 01 0.874 740 702 90.87474070290.874\,740\,702\,90.874 740 702 9 0.283 480 515 20.28348051520.283\,480\,515\,20.283 480 515 2 17.837 395 6117.8373956117.837\,395\,6117.837 395 61 0.726 974 393 50.72697439350.726\,974\,393\,50.726 974 393 5 0.381 584 229 60.38158422960.381\,584\,229\,60.381 584 229 6 15.693 889 815.693889815.693\,889\,815.693 889 8 0.720 088 558 70.72008855870.720\,088\,558\,70.720 088 558 7 0.563 953 320 70.56395332070.563\,953\,320\,70.563 953 320 7 15.873 767 2115.8737672115.873\,767\,2115.873 767 21 0.768 173 2250.7681732250.768\,173\,2250.768 173 225 0.402 816 397 20.40281639720.402\,816\,397\,20.402 816 397 2
Blurry RGB + FreeNeRF [61] 24.877 242 5524.8772425524.877\,242\,5524.877 242 55 0.906 498 842 80.90649884280.906\,498\,842\,80.906 498 842 8 0.118 849 081 10.11884908110.118\,849\,081\,10.118 849 081 1 29.657 663 8929.6576638929.657\,663\,8929.657 663 89 0.941 888 366 50.94188836650.941\,888\,366\,50.941 888 366 5 0.079 584 860 510.079584860510.079\,584\,860\,510.079 584 860 51 19.749 714 8219.7497148219.749\,714\,8219.749 714 82 0.874 473 930.874473930.874\,473\,930.874 473 93 0.222 323 1640.2223231640.222\,323\,1640.222 323 164 20.056 505 5820.0565055820.056\,505\,5820.056 505 58 0.755 278 954 90.75527895490.755\,278\,954\,90.755 278 954 9 0.319 858 026 60.31985802660.319\,858\,026\,60.319 858 026 6 22.277 594 5122.2775945122.277\,594\,5122.277 594 51 0.829 269 746 80.82926974680.829\,269\,746\,80.829 269 746 8 0.167 236 993 50.16723699350.167\,236\,993\,50.167 236 993 5 23.323 744 2723.3237442723.323\,744\,2723.323 744 27 0.861 481 968 20.86148196820.861\,481\,968\,20.861 481 968 2 0.181 570 425 20.18157042520.181\,570\,425\,20.181 570 425 2
GT RGB [37] + Dyn-NeRF $25.683$ $0.912$ $0.159$ $30.234$ $0.944$ $0.106$ $28.132$ $0.915$ $0.107$ $22.088$ $0.794$ $0.312$ $22.840$ $0.819$ $0.228$ $25.795$ $0.877$ $0.182$
GT RGB [37] + NR-NeRF [52] 19.987 609 460 610 0619.9876094606100619.987\,609\,460\,610\,0619.987 609 460 610 06 0.836 801 776 720 3160.8368017767203160.836\,801\,776\,720\,3160.836 801 776 720 316 0.394 049 444 214 345 60.39404944421434560.394\,049\,444\,214\,345\,60.394 049 444 214 345 6 28.789 877 356 542 23828.78987735654223828.789\,877\,356\,542\,23828.789 877 356 542 238 0.934 394 356 739 577 60.93439435673957760.934\,394\,356\,739\,577\,60.934 394 356 739 577 6 0.149 142 221 649 094 380.149142221649094380.149\,142\,221\,649\,094\,380.149 142 221 649 094 38 16.415 376 865 16516.41537686516516.415\,376\,865\,16516.415 376 865 165 0.842 456 244 864 120.842456244864120.842\,456\,244\,864\,120.842 456 244 864 12 0.377 088 615 406 4640.3770886154064640.377\,088\,615\,406\,4640.377 088 615 406 464 14.742 602 446 641 10814.74260244664110814.742\,602\,446\,641\,10814.742 602 446 641 108 0.685 559 039 389 844 10.68555903938984410.685\,559\,039\,389\,844\,10.685 559 039 389 844 1 0.602 584 305 081 0370.6025843050810370.602\,584\,305\,081\,0370.602 584 305 081 037 14.527 485 245 416 34714.52748524541634714.527\,485\,245\,416\,34714.527 485 245 416 347 0.660 709 662 089 724 40.66070966208972440.660\,709\,662\,089\,724\,40.660 709 662 089 724 4 0.580 963 895 659 242 10.58096389565924210.580\,963\,895\,659\,242\,10.580 963 895 659 242 1 18.892 590 274 874 918.892590274874918.892\,590\,274\,874\,918.892 590 274 874 9 0.791 984 215 960 7160.7919842159607160.791\,984\,215\,960\,7160.791 984 215 960 716 0.420 765 696 402 0370.4207656964020370.420\,765\,696\,402\,0370.420 765 696 402 037
Dancing (real) Bucket (real) Towel Tricks (real) Average (real)
Method ROI PSNR $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ROI PSNR $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ROI PSNR $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ ROI PSNR $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Our Dyn-EventNeRF $22.25$ 28.575 158 369 880 528.575158369880528.575\,158\,369\,880\,528.575 158 369 880 5 0.814 792 092 848 1260.8147920928481260.814\,792\,092\,848\,1260.814 792 092 848 126 0.136 793 960 183 8590.1367939601838590.136\,793\,960\,183\,8590.136 793 960 183 859 $24.45$ 29.562 390 701 000 429.562390701000429.562\,390\,701\,000\,429.562 390 701 000 4 0.825 172 266 504 3580.8251722665043580.825\,172\,266\,504\,3580.825 172 266 504 358 0.117 926 983 237 2670.1179269832372670.117\,926\,983\,237\,2670.117 926 983 237 267 $20.82$ 27.597 715 647 052 927.597715647052927.597\,715\,647\,052\,927.597 715 647 052 9 0.808 370 518 658 90.80837051865890.808\,370\,518\,658\,90.808 370 518 658 9 0.135 122 140 571 4750.1351221405714750.135\,122\,140\,571\,4750.135 122 140 571 475 $22.51$ 28.578 421 572 644 628.578421572644628.578\,421\,572\,644\,628.578 421 572 644 6 0.816 111 626 003 7950.8161116260037950.816\,111\,626\,003\,7950.816 111 626 003 795 0.129 947 694 664 20.12994769466420.129\,947\,694\,664\,20.129 947 694 664 2
E2VID [46] + Dyn-NeRF $15.64$ 21.200 901 377 753 121.200901377753121.200\,901\,377\,753\,121.200 901 377 753 1 0.704 781 823 646 8360.7047818236468360.704\,781\,823\,646\,8360.704 781 823 646 836 0.387 421 740 293 5030.3874217402935030.387\,421\,740\,293\,5030.387 421 740 293 503 $16.42$ 18.397 805 267 333 218.397805267333218.397\,805\,267\,333\,218.397 805 267 333 2 0.661 475 154 382 7040.6614751543827040.661\,475\,154\,382\,7040.661 475 154 382 704 0.451 675 924 360 7520.4516759243607520.451\,675\,924\,360\,7520.451 675 924 360 752 $16.71$ 22.780 303 642 668 822.780303642668822.780\,303\,642\,668\,822.780 303 642 668 8 0.732 067 838 703 5740.7320678387035740.732\,067\,838\,703\,5740.732 067 838 703 574 0.307 592 697 441 5780.3075926974415780.307\,592\,697\,441\,5780.307 592 697 441 578 $16.26$ 20.793 003 429 251 720.793003429251720.793\,003\,429\,251\,720.793 003 429 251 7 0.699 441 605 577 7050.6994416055777050.699\,441\,605\,577\,7050.699 441 605 577 705 0.382 230 120 698 6110.3822301206986110.382\,230\,120\,698\,6110.382 230 120 698 611
Blurry RGB + Dyn-NeRF $18.88$ 26.021 426 243 369 826.021426243369826.021\,426\,243\,369\,826.021 426 243 369 8 0.787 472 315 151 3040.7874723151513040.787\,472\,315\,151\,3040.787 472 315 151 304 0.181 774 728 000 1640.1817747280001640.181\,774\,728\,000\,1640.181 774 728 000 164 $21.58$ 28.084 769 859 135 628.084769859135628.084\,769\,859\,135\,628.084 769 859 135 6 0.810 105 238 127 9480.8101052381279480.810\,105\,238\,127\,9480.810 105 238 127 948 0.135 202 926 620 8410.1352029266208410.135\,202\,926\,620\,8410.135 202 926 620 841 $15.42$ 22.059 835 072 862 122.059835072862122.059\,835\,072\,862\,122.059 835 072 862 1 0.735 972 632 006 7540.7359726320067540.735\,972\,632\,006\,7540.735 972 632 006 754 0.246 245 318 502 1880.2462453185021880.246\,245\,318\,502\,1880.246 245 318 502 188 $18.63$ 25.388 677 058 455 825.388677058455825.388\,677\,058\,455\,825.388 677 058 455 8 0.777 850 061 762 0020.7778500617620020.777\,850\,061\,762\,0020.777 850 061 762 002 0.187 740 991 041 0640.1877409910410640.187\,740\,991\,041\,0640.187 740 991 041 064
E2VID [46] + FreeNeRF [61] $10.99$ 9.345 662 947 350 159.345662947350159.345\,662\,947\,350\,159.345 662 947 350 15 0.234 185 938 471 0910.2341859384710910.234\,185\,938\,471\,0910.234 185 938 471 091 1.061 991 870 403 291.061991870403291.061\,991\,870\,403\,291.061 991 870 403 29 $9.72$ 8.997 353 770 071 478.997353770071478.997\,353\,770\,071\,478.997 353 770 071 47 0.221 940 116 537 0960.2219401165370960.221\,940\,116\,537\,0960.221 940 116 537 096 1.033 688 037 168 411.033688037168411.033\,688\,037\,168\,411.033 688 037 168 41 $14.67$ 13.093 696 152 952 913.093696152952913.093\,696\,152\,952\,913.093 696 152 952 9 0.259 799 485 800 8920.2597994858008920.259\,799\,485\,800\,8920.259 799 485 800 892 1.095 114 636 421 21.09511463642121.095\,114\,636\,421\,21.095 114 636 421 2 $11.79$ 10.478 904 290 124 810.478904290124810.478\,904\,290\,124\,810.478 904 290 124 8 0.238 641 846 936 360.238641846936360.238\,641\,846\,936\,360.238 641 846 936 36 1.063 598 181 330 971.063598181330971.063\,598\,181\,330\,971.063 598 181 330 97
EDI [37] + Dyn-NeRF 19.407 062 619 026 35519.40706261902635519.407\,062\,619\,026\,35519.407 062 619 026 355 26.734 226 205 441 6326.7342262054416326.734\,226\,205\,441\,6326.734 226 205 441 63 0.795 924 807 186 280.795924807186280.795\,924\,807\,186\,280.795 924 807 186 28 0.156 415 476 575 493 820.156415476575493820.156\,415\,476\,575\,493\,820.156 415 476 575 493 82 21.851 387 389 761 9721.8513873897619721.851\,387\,389\,761\,9721.851 387 389 761 97 28.312 048 842 436 3528.3120488424363528.312\,048\,842\,436\,3528.312 048 842 436 35 0.811 441 091 722 657 40.81144109172265740.811\,441\,091\,722\,657\,40.811 441 091 722 657 4 0.135 174 065 902 829 180.135174065902829180.135\,174\,065\,902\,829\,180.135 174 065 902 829 18 18.012 607 941 002 0118.0126079410020118.012\,607\,941\,002\,0118.012 607 941 002 01 25.532 457 205 697 28825.53245720569728825.532\,457\,205\,697\,28825.532 457 205 697 288 0.788 793 035 482 977 30.78879303548297730.788\,793\,035\,482\,977\,30.788 793 035 482 977 3 0.154 525 824 636 220 940.154525824636220940.154\,525\,824\,636\,220\,940.154 525 824 636 220 94 19.757 019 316 596 819.757019316596819.757\,019\,316\,596\,819.757 019 316 596 8 26.859 577 417 858 426.859577417858426.859\,577\,417\,858\,426.859 577 417 858 4 0.798 719 644 797 3050.7987196447973050.798\,719\,644\,797\,3050.798 719 644 797 305 0.148 705 122 371 5150.1487051223715150.148\,705\,122\,371\,5150.148 705 122 371 515
EDI [37] + NR-NeRF [52] 14.986 377 826 056 08814.98637782605608814.986\,377\,826\,056\,08814.986 377 826 056 088 15.487 755 160 777 87115.48775516077787115.487\,755\,160\,777\,87115.487 755 160 777 871 0.241 756 347 089 116 970.241756347089116970.241\,756\,347\,089\,116\,970.241 756 347 089 116 97 0.956 517 256 796 360.956517256796360.956\,517\,256\,796\,360.956 517 256 796 36 16.598 653 829 189 09316.59865382918909316.598\,653\,829\,189\,09316.598 653 829 189 093 15.100 904 103 838 80215.10090410383880215.100\,904\,103\,838\,80215.100 904 103 838 802 0.295 238 913 563 5570.2952389135635570.295\,238\,913\,563\,5570.295 238 913 563 557 1.040 737 645 966 666 21.04073764596666621.040\,737\,645\,966\,666\,21.040 737 645 966 666 2 15.353 822 247 858 81915.35382224785881915.353\,822\,247\,858\,81915.353 822 247 858 819 15.816 823 033 775 52215.81682303377552215.816\,823\,033\,775\,52215.816 823 033 775 522 0.239 233 887 558 952 970.239233887558952970.239\,233\,887\,558\,952\,970.239 233 887 558 952 97 1.048 303 231 596 946 71.04830323159694671.048\,303\,231\,596\,946\,71.048 303 231 596 946 7 15.646 284 634 36815.64628463436815.646\,284\,634\,36815.646 284 634 368 15.468 494 099 464 115.468494099464115.468\,494\,099\,464\,115.468 494 099 464 1 0.258 743 049 403 8760.2587430494038760.258\,743\,049\,403\,8760.258 743 049 403 876 1.015 186 044 786 661.015186044786661.015\,186\,044\,786\,661.015 186 044 786 66
Table 1: We compare our method to the baselines on both the synthetic (top) and real (bottom) datasets by rendering novel views and computing scores against the reference images for those views. Dyn-NeRF refers to the proposed Dyn-EventNeRF method with only RGB supervision. Our method produces higher output quality than other methods, consistently across all scenes. ROI scores were computed using a tight per-frame foreground mask estimated from the ground truth. GT RGB is using ground-truth non-blurry training views as input. Despite that, NR-NeRF cannot produce accurate results as the motion is too fast and large for it to learn deformations. Zoom recommended.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours (TensoRF-CP [7]) 24.167 333 584 754 58724.16733358475458724.167\,333\,584\,754\,58724.167 333 584 754 587 0.839 166 336 633 191 80.83916633663319180.839\,166\,336\,633\,191\,80.839 166 336 633 191 8 0.285 319 627 374 410 630.285319627374410630.285\,319\,627\,374\,410\,630.285 319 627 374 410 63
Ours (NGP [31]) 23.239 422 544 419 4223.2394225444194223.239\,422\,544\,419\,4223.239 422 544 419 42 0.776 564 987 385 244 30.77656498738524430.776\,564\,987\,385\,244\,30.776 564 987 385 244 3 0.327 409 669 334 689 730.327409669334689730.327\,409\,669\,334\,689\,730.327 409 669 334 689 73
Ours (HexPlane [4]) 21.694 189 476 829 2521.6941894768292521.694\,189\,476\,829\,2521.694 189 476 829 25 0.803 876 806 325 1890.8038768063251890.803\,876\,806\,325\,1890.803 876 806 325 189 0.363 865 283 777 316 40.36386528377731640.363\,865\,283\,777\,316\,40.363 865 283 777 316 4
w/o clipping 24.470 577 445 092 77224.47057744509277224.470\,577\,445\,092\,77224.470 577 445 092 772 0.864 222 437 643 4620.8642224376434620.864\,222\,437\,643\,4620.864 222 437 643 462 0.240 079 302 365 581 170.240079302365581170.240\,079\,302\,365\,581\,170.240 079 302 365 581 17
w/o $L_{\mathrm{event}}$ 24.510 529 010 252 2924.5105290102522924.510\,529\,010\,252\,2924.510 529 010 252 29 0.863 816 167 892 699 40.86381616789269940.863\,816\,167\,892\,699\,40.863 816 167 892 699 4 0.235 803 375 517 328 60.23580337551732860.235\,803\,375\,517\,328\,60.235 803 375 517 328 6
w/o $L_{\mathrm{acc}}$ 25.113 050 803 055 4525.1130508030554525.113\,050\,803\,055\,4525.113 050 803 055 45 0.859 877 538 668 510 70.85987753866851070.859\,877\,538\,668\,510\,70.859 877 538 668 510 7 0.204 149 595 623 215 050.204149595623215050.204\,149\,595\,623\,215\,050.204 149 595 623 215 05
w/o $L_{\mathrm{RGB}}$ 26.593 070 219 085 88826.59307021908588826.593\,070\,219\,085\,88826.593 070 219 085 888 0.885 689 965 567 769 20.88568996556776920.885\,689\,965\,567\,769\,20.885 689 965 567 769 2 0.158 882 698 702 315 470.158882698702315470.158\,882\,698\,702\,315\,470.158 882 698 702 315 47
only $L_{\mathrm{event}}$ 25.525 104 707 466 88825.52510470746688825.525\,104\,707\,466\,88825.525 104 707 466 888 0.864 045 117 673 428 60.86404511767342860.864\,045\,117\,673\,428\,60.864 045 117 673 428 6 0.202 158 188 459 773 850.202158188459773850.202\,158\,188\,459\,773\,850.202 158 188 459 773 85
only $L_{\mathrm{acc}}$ 25.795 365 393 175 5325.7953653931755325.795\,365\,393\,175\,5325.795 365 393 175 53 0.876 945 422 335 5340.8769454223355340.876\,945\,422\,335\,5340.876 945 422 335 534 0.182 430 626 079 440 130.182430626079440130.182\,430\,626\,079\,440\,130.182 430 626 079 440 13
w/o $L_{\mathrm{sparsity}}$ 26.853 168 140 108 5626.8531681401085626.853\,168\,140\,108\,5626.853 168 140 108 56 0.886 839 200 463 5490.8868392004635490.886\,839\,200\,463\,5490.886 839 200 463 549 0.157 080 707 655 598 730.157080707655598730.157\,080\,707\,655\,598\,730.157 080 707 655 598 73
Our Full Model (Final) 27.096 711 660 788 7527.0967116607887527.096\,711\,660\,788\,7527.096 711 660 788 75 0.893 037 778 571 1420.8930377785711420.893\,037\,778\,571\,1420.893 037 778 571 142 0.145 413 492 266 088 70.14541349226608870.145\,413\,492\,266\,088\,70.145 413 492 266 088 7
Table 2: Quantitative ablation and design choice study averaged over all synthetic scenes. All metrics clearly favour our full model.
5.3 Implementation
[…]
5.4 Comparisons
[…]
Our Dynamic EventNeRF (abbr.: Dyn-EventNeRF) outperforms all of the baselines (Tab. 1). Figs. 3 and 2 confirm this visually: All baselines based on blurry RGBs fail to recover sharp details as they were designed to work with sharp data. In contrast, our proposed Dyn-EventNeRF method uses events instead, which allows it to recover sharp details. Methods trained on RGB frames generated from events using E2VID and EDI also recover these details. However, similarly to Ref. [49], there are many artefacts in the reconstructions, as all input views were reconstructed independently, disregarding multi-view consistency. Our method integrates multi-view events into one shared volume, allowing for view-consistent 3D reconstruction and reduced artefacts. NR-NeRF [52] fails to reconstruct most of the scenes. Due to the extreme amplitude and speed of the motion and view sparsity in the tested sequences, it can learn neither the deformations nor the canonical volume correctly.
5.5 Ablation and Design Choice Study
We conduct an ablation and design choice study (Tabs. 2, 5 and 4; per-scene in Tab. 4 in the appendix) using both synthetic and real data. In particular, we compare different options for the core model: MLP (“Full Model”), NGP [31], TensoRF-CP [7], and HexPlane [4]. NGP fails to handle the sparse-view setting. TensoRF-CP and HexPlane are both grid-based: Individual timestamps are modelled separately, so that information from one timestamp cannot propagate to other timestamps, leading to blurriness and artefacts. On the other hand, our full model with temporally-conditioned MLP explicitly controls how much information is shared between timestamps through positional encoding. Supervising one timestamp also affects others in its vicinity, resulting in better use of the training data. Disabling $\mathcal{L}_{\text{event}}$, $\mathcal{L}_{\text{acc}}$, or $\mathcal{L}_{\text{RGB}}$ leads to increased blurriness because disabling either of them will reduce the amount of supervision: Without $\mathcal{L}_{\text{event}}$, there is no way to correct relative changes between predictions through events, without $\mathcal{L}_{\text{acc}}$ there is no way to propagate sharp RGB information and static parts of the scene. And without $\mathcal{L}_{\text{RGB}}$, the model is not encouraged to predict the reference RGB frames correctly. Disabling both $\mathcal{L}_{\text{RGB}}$ and $\mathcal{L}_{\text{acc}}$ removes the entire RGB supervision. As the model only uses the events, there can be an exposure ambiguity, leading to lower metrics. Disabling both $\mathcal{L}_{\text{RGB}}$ and $\mathcal{L}_{\text{event}}$ leaves only $\mathcal{L}_{\text{acc}}$. Despite using both RGB and events, this baseline performs worse than the full model, presumably due to accumulation artefacts. Cylinder clipping prevents the model from creating geometry in regions that are only supervised by one view. If disabled, the model produces numerous artefacts in the novel views; in the case of real data specifically, it even fails to converge. $\mathcal{L}_{\text{sparsity}}$ forces the model to reduce the number of opaque pixels in the reconstruction. Removing it leads to semi-transparent floating artefacts in the reconstructed volume. Using only one single model for the entire sequence duration (as opposed to dividing the sequence into shorter chunks) results in a blurrier reconstruction. In this case, artefacts produced by reconstructing one of the segments can propagate to other segments, which is not the case in the multi-segment model.
6 Conclusion
We introduced Dynamic EventNeRF—the first approach to 4D reconstruction of dynamic, non-rigid scenes from multi-view event streams augmented with sparse RGB frames. Our approach outperforms all tested baselines, producing state-of-the-art reconstructions of fast motions under dim and dark lighting conditions. Our ablation study shows that our full model, comprising events, RGB, sparsity, event accumulation losses and cylinder clipping performs best. In particular, we found the MLP backbone to be the most suitable for the task.
In a further qualitative ablation study (see Appx. G), we have analysed the impact of the number of viewpoints on the reconstruction quality. The experiment indicates that quality is indeed increased by adding additional views. This can serve as a justification for further research into multi-view event camera setups for 4D reconstruction.
Adding external high-resolution RGB cameras to the event setup could further improve the sharpness of the rendered novel views. Future work could also investigate architectures that achieve similar quality at a smaller cost (e.g. a variant of 3DGS [17]).
We assembled and calibrated a multi-view event camera setup to record, to the best of our knowledge, the first dataset that consists of six-view simultaneous events and RGB recordings of complex motions in challenging lighting conditions. We made this dataset available online, such that despite the limited availability of multi-view event camera data, further research can continue to explore this modality.
Appendices
[…]
Figure 7: RGB frames recorded by the DAVIS camera are very blurry, because low-light conditions require longer exposure times.
This appendix provides additional experiments and details on calibration, hyperparameters and baselines. We show how we calibrate our event camera response function in Appx. B. In Appx. C, we describe the architecture of our MLP network. Next, Appx. D includes additional details on the baselines and their hyperparameters. We then describe our datasets and explain the capture process in Appx. E. We demonstrate the performance of our method with additional ablations on the number of supporting RGB frames (Appx. F), number of views (Appx. G), and provide the full per-scene ablation results in Tab. 4. Finally, we discuss event accumulation stability and the proposed accumulation decay, proving it effective and correct in our case (Appx. I).
(a) Raw recorded RGB values when varying the exposure time of the camera at different pixel locations and averaged over the colour channels. There is an outlier on the right of the blue channel average curve, which we ignore during calibration.
[…]
Figure 8: Event camera RGB intensity frame CRF calibration.
Appendix A RGB Frame Quality
As a result of the low-light conditions (see Sec. 5.2), the exposure durations of the RGB frames recorded by the DAVIS camera are rather long, resulting in severe motion blur (see Fig. 7).
Appendix B Camera CRF Calibration
[…]
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours (TensoRF-CP [7]) 26.201 907 4526.2019074526.201\,907\,4526.201 907 45 0.784 033 3810.7840333810.784\,033\,3810.784 033 381 0.163 783 0910.1637830910.163\,783\,0910.163 783 091
Ours (NGP [31]) 26.278 212 4426.2782124426.278\,212\,4426.278 212 44 0.787 437 7110.7874377110.787\,437\,7110.787 437 711 0.163 224 6900.1632246900.163\,224\,6900.163 224 690
Ours (HexPlane [4]) 24.940 644 7524.9406447524.940\,644\,7524.940 644 75 0.758 267 7500.7582677500.758\,267\,7500.758 267 750 0.209 537 8970.2095378970.209\,537\,8970.209 537 897
w/o clipping 25.287 940 2925.2879402925.287\,940\,2925.287 940 29 0.809 762 0550.8097620550.809\,762\,0550.809 762 055 0.162 667 8270.1626678270.162\,667\,8270.162 667 827
w/o decay 27.118 974 8827.1189748827.118\,974\,8827.118 974 88 0.815 725 3050.8157253050.815\,725\,3050.815 725 305 0.133 418 8450.1334188450.133\,418\,8450.133 418 845
w/o multi-segment 26.672 972 9226.6729729226.672\,972\,9226.672 972 92 0.809 257 5750.8092575750.809\,257\,5750.809 257 575 0.142 403 3890.1424033890.142\,403\,3890.142 403 389
w/o $L_{\mathrm{sparsity}}$ 26.919 654 8526.9196548526.919\,654\,8526.919 654 85 0.813 790 6260.8137906260.813\,790\,6260.813 790 626 0.127 648 8630.1276488630.127\,648\,8630.127 648 863
w/o $L_{\mathrm{event}}$ 27.619 807 4627.6198074627.619\,807\,4627.619 807 46 0.820 358 3380.8203583380.820\,358\,3380.820 358 338 0.122 953 4890.1229534890.122\,953\,4890.122 953 489
w/o $L_{\mathrm{acc}}$ 25.517 277 1625.5172771625.517\,277\,1625.517 277 16 0.801 984 9870.8019849870.801\,984\,9870.801 984 987 0.139 127 6000.1391276000.139\,127\,6000.139 127 600
w/o $L_{\mathrm{RGB}}$ 27.061 607 1327.0616071327.061\,607\,1327.061 607 13 0.818 148 0330.8181480330.818\,148\,0330.818 148 033 0.120 109 7500.1201097500.120\,109\,7500.120 109 750
only $L_{\mathrm{event}}$ 25.028 554 8725.0285548725.028\,554\,8725.028 554 87 0.798 512 3700.7985123700.798\,512\,3700.798 512 370 0.142 937 2800.1429372800.142\,937\,2800.142 937 280
only $L_{\mathrm{acc}}$ 27.754 136 9327.7541369327.754\,136\,9327.754 136 93 0.821 195 0910.8211950910.821\,195\,0910.821 195 091 0.121 601 9530.1216019530.121\,601\,9530.121 601 953
only $L_{\mathrm{RGB}}$ 26.019 370 863 660 69426.01937086366069426.019\,370\,863\,660\,69426.019 370 863 660 694 0.805 922 053 215 070 70.80592205321507070.805\,922\,053\,215\,070\,70.805 922 053 215 070 7 0.144 494 439 586 997 040.144494439586997040.144\,494\,439\,586\,997\,040.144 494 439 586 997 04
Our Full Model (Final) 27.047 641 4027.0476414027.047\,641\,4027.047 641 40 0.818 727 1740.8187271740.818\,727\,1740.818 727 174 0.119 746 3490.1197463490.119\,746\,3490.119 746 349
Table 3: Quantitative ablation and design choice study on the “Sword” real scene. SSIM and LPIPS metrics favour our full model. In particular, omitting event decay (”w/o decay”) makes SSIM and LPIPS scores worse.
Blender Dress Spheres
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
TensoRF-CP [7] 24.726 813 089 903 86424.72681308990386424.726\,813\,089\,903\,86424.726 813 089 903 864 0.879 434 371 921 662 70.87943437192166270.879\,434\,371\,921\,662\,70.879 434 371 921 662 7 0.226 776 018 987 099 340.226776018987099340.226\,776\,018\,987\,099\,340.226 776 018 987 099 34 28.090 594 246 819 5228.0905942468195228.090\,594\,246\,819\,5228.090 594 246 819 52 0.914 956 114 329 549 20.91495611432954920.914\,956\,114\,329\,549\,20.914 956 114 329 549 2 0.202 168 937 772 512 440.202168937772512440.202\,168\,937\,772\,512\,440.202 168 937 772 512 44 23.971 144 235 363 90523.97114423536390523.971\,144\,235\,363\,90523.971 144 235 363 905 0.866 284 227 226 692 40.86628422722669240.866\,284\,227\,226\,692\,40.866 284 227 226 692 4 0.279 547 073 692 083 370.279547073692083370.279\,547\,073\,692\,083\,370.279 547 073 692 083 37
NGP [31] 25.687 043 317 672 54325.68704331767254325.687\,043\,317\,672\,54325.687 043 317 672 543 0.888 347 364 472 563 10.88834736447256310.888\,347\,364\,472\,563\,10.888 347 364 472 563 1 0.184 152 255 704 005 560.184152255704005560.184\,152\,255\,704\,005\,560.184 152 255 704 005 56 29.131 450 056 670 45629.13145005667045629.131\,450\,056\,670\,45629.131 450 056 670 456 0.933 399 358 625 454 10.93339935862545410.933\,399\,358\,625\,454\,10.933 399 358 625 454 1 0.178 910 806 402 564 040.178910806402564040.178\,910\,806\,402\,564\,040.178 910 806 402 564 04 28.754 522 901 386 3928.7545229013863928.754\,522\,901\,386\,3928.754 522 901 386 39 0.914 640 666 232 544 60.91464066623254460.914\,640\,666\,232\,544\,60.914 640 666 232 544 6 0.120 185 600 221 157 070.120185600221157070.120\,185\,600\,221\,157\,070.120 185 600 221 157 07
HexPlane [4] 23.119 187 666 592 20723.11918766659220723.119\,187\,666\,592\,20723.119 187 666 592 207 0.863 639 933 196 624 40.86363993319662440.863\,639\,933\,196\,624\,40.863 639 933 196 624 4 0.274 748 245 378 335 30.27474824537833530.274\,748\,245\,378\,335\,30.274 748 245 378 335 3 23.580 262 647 709 40423.58026264770940423.580\,262\,647\,709\,40423.580 262 647 709 404 0.866 452 846 424 437 50.86645284642443750.866\,452\,846\,424\,437\,50.866 452 846 424 437 5 0.335 567 439 844 210 950.335567439844210950.335\,567\,439\,844\,210\,950.335 567 439 844 210 95 21.636 131 887 597 45821.63613188759745821.636\,131\,887\,597\,45821.636 131 887 597 458 0.853 130 213 071 383 10.85313021307138310.853\,130\,213\,071\,383\,10.853 130 213 071 383 1 0.333 956 459 164 619 470.333956459164619470.333\,956\,459\,164\,619\,470.333 956 459 164 619 47
w/o clipping 19.958 524 002 730 619.958524002730619.958\,524\,002\,730\,619.958 524 002 730 6 0.851 266 843 374 816 60.85126684337481660.851\,266\,843\,374\,816\,60.851 266 843 374 816 6 0.425 993 429 621 060 70.42599342962106070.425\,993\,429\,621\,060\,70.425 993 429 621 060 7 27.926 075 186 536 85627.92607518653685627.926\,075\,186\,536\,85627.926 075 186 536 856 0.925 761 656 247 033 80.92576165624703380.925\,761\,656\,247\,033\,80.925 761 656 247 033 8 0.146 590 058 195 094 270.146590058195094270.146\,590\,058\,195\,094\,270.146 590 058 195 094 27 28.785 698 739 645 66428.78569873964566428.785\,698\,739\,645\,66428.785 698 739 645 664 0.916 252 704 153 087 50.91625270415308750.916\,252\,704\,153\,087\,50.916 252 704 153 087 5 0.093 198 148 844 142 750.093198148844142750.093\,198\,148\,844\,142\,750.093 198 148 844 142 75
w/o $L_{\mathrm{sparsity}}$ 27.500 890 010 806 58627.50089001080658627.500\,890\,010\,806\,58627.500 890 010 806 586 0.926 403 250 650 059 40.92640325065005940.926\,403\,250\,650\,059\,40.926 403 250 650 059 4 0.134 559 555 413 822 330.134559555413822330.134\,559\,555\,413\,822\,330.134 559 555 413 822 33 31.292 436 937 216 57531.29243693721657531.292\,436\,937\,216\,57531.292 436 937 216 575 0.950 075 811 437 750 80.95007581143775080.950\,075\,811\,437\,750\,80.950 075 811 437 750 8 0.081 498 009 028 534 090.081498009028534090.081\,498\,009\,028\,534\,090.081 498 009 028 534 09 29.357 744 778 639 54229.35774477863954229.357\,744\,778\,639\,54229.357 744 778 639 542 0.920 189 548 561 450.920189548561450.920\,189\,548\,561\,450.920 189 548 561 45 0.085 555 453 008 661 670.085555453008661670.085\,555\,453\,008\,661\,670.085 555 453 008 661 67
w/o $L_{\mathrm{event}}$ 19.958 524 002 730 619.958524002730619.958\,524\,002\,730\,619.958 524 002 730 6 0.851 266 843 374 816 60.85126684337481660.851\,266\,843\,374\,816\,60.851 266 843 374 816 6 0.425 993 429 621 060 70.42599342962106070.425\,993\,429\,621\,060\,70.425 993 429 621 060 7 29.883 649 189 577 00329.88364918957700329.883\,649\,189\,577\,00329.883 649 189 577 003 0.941 468 641 695 164 40.94146864169516440.941\,468\,641\,695\,164\,40.941 468 641 695 164 4 0.108 069 653 188 188 870.108069653188188870.108\,069\,653\,188\,188\,870.108 069 653 188 188 87 27.744 175 527 732 09327.74417552773209327.744\,175\,527\,732\,09327.744 175 527 732 093 0.913 152 292 681 087 80.91315229268108780.913\,152\,292\,681\,087\,80.913 152 292 681 087 8 0.113 817 538 072 665 540.113817538072665540.113\,817\,538\,072\,665\,540.113 817 538 072 665 54
w/o $L_{\mathrm{acc}}$ 26.834 836 707 336 91526.83483670733691526.834\,836\,707\,336\,91526.834 836 707 336 915 0.899 462 385 654 854 90.89946238565485490.899\,462\,385\,654\,854\,90.899 462 385 654 854 9 0.124 908 758 327 364 920.124908758327364920.124\,908\,758\,327\,364\,920.124 908 758 327 364 92 30.649 549 771 314 9830.6495497713149830.649\,549\,771\,314\,9830.649 549 771 314 98 0.946 437 680 159 603 10.94643768015960310.946\,437\,680\,159\,603\,10.946 437 680 159 603 1 0.078 871 154 909 332 60.07887115490933260.078\,871\,154\,909\,332\,60.078 871 154 909 332 6 28.846 532 259 455 54328.84653225945554328.846\,532\,259\,455\,54328.846 532 259 455 543 0.919 801 924 609 636 50.91980192460963650.919\,801\,924\,609\,636\,50.919 801 924 609 636 5 0.087 039 245 044 191 680.087039245044191680.087\,039\,245\,044\,191\,680.087 039 245 044 191 68
w/o $L_{\mathrm{RGB}}$ 27.320 703 359 867 5927.3207033598675927.320\,703\,359\,867\,5927.320 703 359 867 59 0.927 934 694 661 1580.9279346946611580.927\,934\,694\,661\,1580.927 934 694 661 158 0.121 917 983 517 050 740.121917983517050740.121\,917\,983\,517\,050\,740.121 917 983 517 050 74 30.540 144 287 499 25630.54014428749925630.540\,144\,287\,499\,25630.540 144 287 499 256 0.945 758 354 501 286 50.94575835450128650.945\,758\,354\,501\,286\,50.945 758 354 501 286 5 0.085 657 446 707 288 420.085657446707288420.085\,657\,446\,707\,288\,420.085 657 446 707 288 42 29.128 532 219 304 41829.12853221930441829.128\,532\,219\,304\,41829.128 532 219 304 418 0.920 345 419 224 526 60.92034541922452660.920\,345\,419\,224\,526\,60.920 345 419 224 526 6 0.088 927 259 234 090 640.088927259234090640.088\,927\,259\,234\,090\,640.088 927 259 234 090 64
only $L_{\mathrm{event}}$ 27.361 913 593 041 24827.36191359304124827.361\,913\,593\,041\,24827.361 913 593 041 248 0.905 326 513 363 621 90.90532651336362190.905\,326\,513\,363\,621\,90.905 326 513 363 621 9 0.147 916 332 632 303 250.147916332632303250.147\,916\,332\,632\,303\,250.147 916 332 632 303 25 31.307 773 242 231 731.307773242231731.307\,773\,242\,231\,731.307 773 242 231 7 0.950 624 138 984 662 90.95062413898466290.950\,624\,138\,984\,662\,90.950 624 138 984 662 9 0.072 787 756 038 208 80.07278775603820880.072\,787\,756\,038\,208\,80.072 787 756 038 208 8 29.446 769 051 923 8629.4467690519238629.446\,769\,051\,923\,8629.446 769 051 923 86 0.920 639 318 208 602 30.92063931820860230.920\,639\,318\,208\,602\,30.920 639 318 208 602 3 0.083 587 149 158 120 160.083587149158120160.083\,587\,149\,158\,120\,160.083 587 149 158 120 16
only $L_{\mathrm{acc}}$ 25.683 241 879 154 76725.68324187915476725.683\,241\,879\,154\,76725.683 241 879 154 767 0.912 295 106 325 397 30.91229510632539730.912\,295\,106\,325\,397\,30.912 295 106 325 397 3 0.158 982 390 537 8580.1589823905378580.158\,982\,390\,537\,8580.158 982 390 537 858 30.234 220 903 611 50430.23422090361150430.234\,220\,903\,611\,50430.234 220 903 611 504 0.943 951 205 382 188 20.94395120538218820.943\,951\,205\,382\,188\,20.943 951 205 382 188 2 0.105 726 395 236 949 130.105726395236949130.105\,726\,395\,236\,949\,130.105 726 395 236 949 13 28.132 232 672 459 7228.1322326724597228.132\,232\,672\,459\,7228.132 232 672 459 72 0.915 299 381 578 3680.9152993815783680.915\,299\,381\,578\,3680.915 299 381 578 368 0.107 350 979 807 476 20.10735097980747620.107\,350\,979\,807\,476\,20.107 350 979 807 476 2
Full Model 27.430 956 492 478 71727.43095649247871727.430\,956\,492\,478\,71727.430 956 492 478 717 0.929 069 216 594 528 70.92906921659452870.929\,069\,216\,594\,528\,70.929 069 216 594 528 7 0.119 141 521 615 286 670.119141521615286670.119\,141\,521\,615\,286\,670.119 141 521 615 286 67 31.395 668 613 445 32431.39566861344532431.395\,668\,613\,445\,32431.395 668 613 445 324 0.951 551 222 201 578 30.95155122220157830.951\,551\,222\,201\,578\,30.951 551 222 201 578 3 0.075 906 729 449 828 470.075906729449828470.075\,906\,729\,449\,828\,470.075 906 729 449 828 47 29.765 333 834 938 9629.7653338349389629.765\,333\,834\,938\,9629.765 333 834 938 96 0.924 179 585 104 031 50.92417958510403150.924\,179\,585\,104\,031\,50.924 179 585 104 031 5 0.080 554 159 637 540 580.080554159637540580.080\,554\,159\,637\,540\,580.080 554 159 637 540 58
Lego Static Lego Average
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
TensoRF-CP [7] 21.444 033 233 458 91421.44403323345891421.444\,033\,233\,458\,91421.444 033 233 458 914 0.751 667 764 708 832 90.75166776470883290.751\,667\,764\,708\,832\,90.751 667 764 708 832 9 0.407 979 355 752 468 130.407979355752468130.407\,979\,355\,752\,468\,130.407 979 355 752 468 13 22.604 083 118 226 72222.60408311822672222.604\,083\,118\,226\,72222.604 083 118 226 722 0.783 489 204 979 222 10.78348920497922210.783\,489\,204\,979\,222\,10.783 489 204 979 222 1 0.310 126 750 667 889 90.31012675066788990.310\,126\,750\,667\,889\,90.310 126 750 667 889 9 24.167 333 584 754 58724.16733358475458724.167\,333\,584\,754\,58724.167 333 584 754 587 0.839 166 336 633 191 80.83916633663319180.839\,166\,336\,633\,191\,80.839 166 336 633 191 8 0.285 319 627 374 410 630.285319627374410630.285\,319\,627\,374\,410\,630.285 319 627 374 410 63
NGP [31] 17.134 322 523 579 63417.13432252357963417.134\,322\,523\,579\,63417.134 322 523 579 634 0.594 455 333 820 9390.5944553338209390.594\,455\,333\,820\,9390.594 455 333 820 939 0.556 270 637 114 842 70.55627063711484270.556\,270\,637\,114\,842\,70.556 270 637 114 842 7 15.489 773 922 788 0815.4897739227880815.489\,773\,922\,788\,0815.489 773 922 788 08 0.551 982 213 774 720 80.55198221377472080.551\,982\,213\,774\,720\,80.551 982 213 774 720 8 0.597 529 047 230 879 50.59752904723087950.597\,529\,047\,230\,879\,50.597 529 047 230 879 5 23.239 422 544 419 4223.2394225444194223.239\,422\,544\,419\,4223.239 422 544 419 42 0.776 564 987 385 244 30.77656498738524430.776\,564\,987\,385\,244\,30.776 564 987 385 244 3 0.327 409 669 334 689 730.327409669334689730.327\,409\,669\,334\,689\,730.327 409 669 334 689 73
HexPlane [4] 20.093 183 911 957 4420.0931839119574420.093\,183\,911\,957\,4420.093 183 911 957 44 0.711 099 451 852 795 90.71109945185279590.711\,099\,451\,852\,795\,90.711 099 451 852 795 9 0.451 924 413 442 611 70.45192441344261170.451\,924\,413\,442\,611\,70.451 924 413 442 611 7 20.042 181 270 289 74520.04218127028974520.042\,181\,270\,289\,74520.042 181 270 289 745 0.725 061 587 080 704 10.72506158708070410.725\,061\,587\,080\,704\,10.725 061 587 080 704 1 0.423 129 861 056 804 670.423129861056804670.423\,129\,861\,056\,804\,670.423 129 861 056 804 67 21.694 189 476 829 2521.6941894768292521.694\,189\,476\,829\,2521.694 189 476 829 25 0.803 876 806 325 1890.8038768063251890.803\,876\,806\,325\,1890.803 876 806 325 189 0.363 865 283 777 316 40.36386528377731640.363\,865\,283\,777\,316\,40.363 865 283 777 316 4
w/o clipping 22.995 822 244 615 4422.9958222446154422.995\,822\,244\,615\,4422.995 822 244 615 44 0.813 506 704 485 771 20.81350670448577120.813\,506\,704\,485\,771\,20.813 506 704 485 771 2 0.272 479 312 866 926 170.272479312866926170.272\,479\,312\,866\,926\,170.272 479 312 866 926 17 22.686 767 051 935 2922.6867670519352922.686\,767\,051\,935\,2922.686 767 051 935 29 0.814 324 279 956 600 60.81432427995660060.814\,324\,279\,956\,600\,60.814 324 279 956 600 6 0.262 135 562 300 682 10.26213556230068210.262\,135\,562\,300\,682\,10.262 135 562 300 682 1 24.470 577 445 092 77224.47057744509277224.470\,577\,445\,092\,77224.470 577 445 092 772 0.864 222 437 643 4620.8642224376434620.864\,222\,437\,643\,4620.864 222 437 643 462 0.240 079 302 365 581 170.240079302365581170.240\,079\,302\,365\,581\,170.240 079 302 365 581 17
w/o $L_{\mathrm{sparsity}}$ 22.416 481 922 556 21622.41648192255621622.416\,481\,922\,556\,21622.416 481 922 556 216 0.798 495 934 024 220 50.79849593402422050.798\,495\,934\,024\,220\,50.798 495 934 024 220 5 0.281 137 408 316 135 430.281137408316135430.281\,137\,408\,316\,135\,430.281 137 408 316 135 43 23.698 287 051 323 87423.69828705132387423.698\,287\,051\,323\,87423.698 287 051 323 874 0.839 031 457 644 264 10.83903145764426410.839\,031\,457\,644\,264\,10.839 031 457 644 264 1 0.202 653 112 510 840 10.20265311251084010.202\,653\,112\,510\,840\,10.202 653 112 510 840 1 26.853 168 140 108 5626.8531681401085626.853\,168\,140\,108\,5626.853 168 140 108 56 0.886 839 200 463 5490.8868392004635490.886\,839\,200\,463\,5490.886 839 200 463 549 0.157 080 707 655 598 730.157080707655598730.157\,080\,707\,655\,598\,730.157 080 707 655 598 73
w/o $L_{\mathrm{event}}$ 21.979 894 253 704 1821.9798942537041821.979\,894\,253\,704\,1821.979 894 253 704 18 0.791 976 209 177 054 90.79197620917705490.791\,976\,209\,177\,054\,90.791 976 209 177 054 9 0.305 459 981 908 400 850.305459981908400850.305\,459\,981\,908\,400\,850.305 459 981 908 400 85 22.986 402 077 517 60322.98640207751760322.986\,402\,077\,517\,60322.986 402 077 517 603 0.821 216 852 535 374 20.82121685253537420.821\,216\,852\,535\,374\,20.821 216 852 535 374 2 0.225 676 274 796 326 960.225676274796326960.225\,676\,274\,796\,326\,960.225 676 274 796 326 96 24.510 529 010 252 2924.5105290102522924.510\,529\,010\,252\,2924.510 529 010 252 29 0.863 816 167 892 699 40.86381616789269940.863\,816\,167\,892\,699\,40.863 816 167 892 699 4 0.235 803 375 517 328 60.23580337551732860.235\,803\,375\,517\,328\,60.235 803 375 517 328 6
w/o $L_{\mathrm{acc}}$ 23.177 754 272 854 6323.1777542728546323.177\,754\,272\,854\,6323.177 754 272 854 63 0.820 089 738 618 533 60.82008973861853360.820\,089\,738\,618\,533\,60.820 089 738 618 533 6 0.246 794 022 123 018 90.24679402212301890.246\,794\,022\,123\,018\,90.246 794 022 123 018 9 16.056 581 004 315 216.056581004315216.056\,581\,004\,315\,216.056 581 004 315 2 0.713 595 964 299 925 20.71359596429992520.713\,595\,964\,299\,925\,20.713 595 964 299 925 2 0.483 134 797 712 167 10.48313479771216710.483\,134\,797\,712\,167\,10.483 134 797 712 167 1 25.113 050 803 055 4525.1130508030554525.113\,050\,803\,055\,4525.113 050 803 055 45 0.859 877 538 668 510 70.85987753866851070.859\,877\,538\,668\,510\,70.859 877 538 668 510 7 0.204 149 595 623 215 050.204149595623215050.204\,149\,595\,623\,215\,050.204 149 595 623 215 05
w/o $L_{\mathrm{RGB}}$ 22.331 562 843 643 8622.3315628436438622.331\,562\,843\,643\,8622.331 562 843 643 86 0.798 587 015 164 470 30.79858701516447030.798\,587\,015\,164\,470\,30.798 587 015 164 470 3 0.292 103 720 953 067 140.292103720953067140.292\,103\,720\,953\,067\,140.292 103 720 953 067 14 23.644 408 385 114 31423.64440838511431423.644\,408\,385\,114\,31423.644 408 385 114 314 0.835 824 344 287 404 20.83582434428740420.835\,824\,344\,287\,404\,20.835 824 344 287 404 2 0.205 807 083 100 080 50.20580708310008050.205\,807\,083\,100\,080\,50.205 807 083 100 080 5 26.593 070 219 085 88826.59307021908588826.593\,070\,219\,085\,88826.593 070 219 085 888 0.885 689 965 567 769 20.88568996556776920.885\,689\,965\,567\,769\,20.885 689 965 567 769 2 0.158 882 698 702 315 470.158882698702315470.158\,882\,698\,702\,315\,470.158 882 698 702 315 47
only $L_{\mathrm{event}}$ 23.009 271 055 062 57723.00927105506257723.009\,271\,055\,062\,57723.009 271 055 062 577 0.814 492 223 583 378 50.81449222358337850.814\,492\,223\,583\,378\,50.814 492 223 583 378 5 0.262 380 520 751 078 90.26238052075107890.262\,380\,520\,751\,078\,90.262 380 520 751 078 9 16.499 796 595 075 03616.49979659507503616.499\,796\,595\,075\,03616.499 796 595 075 036 0.729 143 394 226 877 80.72914339422687780.729\,143\,394\,226\,877\,80.729 143 394 226 877 8 0.444 119 183 719 158 150.444119183719158150.444\,119\,183\,719\,158\,150.444 119 183 719 158 15 25.525 104 707 466 88825.52510470746688825.525\,104\,707\,466\,88825.525 104 707 466 888 0.864 045 117 673 428 60.86404511767342860.864\,045\,117\,673\,428\,60.864 045 117 673 428 6 0.202 158 188 459 773 850.202158188459773850.202\,158\,188\,459\,773\,850.202 158 188 459 773 85
only $L_{\mathrm{acc}}$ 22.087 511 700 456 23622.08751170045623622.087\,511\,700\,456\,23622.087 511 700 456 236 0.793 720 400 351 434 10.79372040035143410.793\,720\,400\,351\,434\,10.793 720 400 351 434 1 0.311 890 050 023 794 160.311890050023794160.311\,890\,050\,023\,794\,160.311 890 050 023 794 16 22.839 619 810 195 43722.83961981019543722.839\,619\,810\,195\,43722.839 619 810 195 437 0.819 461 018 040 281 60.81946101804028160.819\,461\,018\,040\,281\,60.819 461 018 040 281 6 0.228 203 314 791 123 070.228203314791123070.228\,203\,314\,791\,123\,070.228 203 314 791 123 07 25.795 365 393 175 5325.7953653931755325.795\,365\,393\,175\,5325.795 365 393 175 53 0.876 945 422 335 5340.8769454223355340.876\,945\,422\,335\,5340.876 945 422 335 534 0.182 430 626 079 440 130.182430626079440130.182\,430\,626\,079\,440\,130.182 430 626 079 440 13
Full Model 23.029 062 230 350 39823.02906223035039823.029\,062\,230\,350\,39823.029 062 230 350 398 0.818 088 664 897 111 40.81808866489711140.818\,088\,664\,897\,111\,40.818 088 664 897 111 4 0.258 308 009 058 237 10.25830800905823710.258\,308\,009\,058\,237\,10.258 308 009 058 237 1 23.862 537 132 730 34223.86253713273034223.862\,537\,132\,730\,34223.862 537 132 730 342 0.842 300 204 058 459 80.84230020405845980.842\,300\,204\,058\,459\,80.842 300 204 058 459 8 0.193 157 041 569 550 820.193157041569550820.193\,157\,041\,569\,550\,820.193 157 041 569 550 82 27.096 711 660 788 7527.0967116607887527.096\,711\,660\,788\,7527.096 711 660 788 75 0.893 037 778 571 1420.8930377785711420.893\,037\,778\,571\,1420.893 037 778 571 142 0.145 413 492 266 088 70.14541349226608870.145\,413\,492\,266\,088\,70.145 413 492 266 088 7
Table 4: Quantitative ablation and design choice study done with all synthetic scenes. While some ablated models performed better in single scenes, the averaged metrics clearly favour our full model.
Appendix C MLP Architecture and Hyperparameters
We inherit the NeRF MLP architecture used in EventNeRF [49], which we also optimise with Adam [18]. However, we modify the model to use the same shared network for both fine and coarse levels of prediction rather than using separate ones. This allows for better optimisation stability and speed, as only half as many parameters are optimised, compared to the original version. Due to the small number of input views in our setting, we modified the code to compose training batches such that they contain an equal number of rays from each view. We found this to increase the stability of the training and the accuracy of the predictions.
Appendix D Baseline Implementation Details
NGP, TensoRF-CP and HexPlane were reimplemented on top of our codebase. For NGP, we used a hash grid implementation in tiny-cuda-nn [30]. For the hash-grid encoding, we used the following configuration:
    "otype": "HashGrid",
    "n_levels": 8,
    "n_features_per_level": 2,
    "log2_hashmap_size": 19,
    "base_resolution": 8,
    "per_level_scale": 2.0
[…]
Appendix E Dataset Composition
The proposed synthetic dataset consists of
  1. 1.
    Three new original scenes: “Spheres”, “Blender”, “Dress” (licensed CC-BY4.0), and
  2. 2.
    Two scenes that were based on the data provided in [26]: “Lego”, “Static Lego”.
[…]
We captured ten subjects. Each of them was instructed about the recording and signed the release form for the use of the recorded data in our experiments and subsequent public release. The instructions were as follows: “Enter the recording area. Run around the centre of the area, perform kicks, punches, jumping jacks, and then whatever fast motions you like for a total of a minute. Afterwards, take one of the available objects (towel, ball, bass guitar, paper poster ‘sword’, box, bucket) and perform fast motions for about another minute.”
0.5 FPS 1 FPS 5 FPS 10 FPS 50 FPS 100 FPS Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 9: Ablation on the number of supporting RGB images used for training. We show novel views from two different times in two rows; bold indicates the default value. The 0.5 FPS model uses only one set of RGB images. As the number of images increases, there is a slight reduction in artefacts. However, even with one set of images (0.5 FPS), the results are close to the full model (5 FPS). This indicates that the method uses primarily event information and does not rely on the RGB images much.
2 views 3 views 4 views 5 views Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: Ablation on the number of views used for training. We show novel views at two different times in two rows; bold indicates the default value. With more input views, the quality indeed significantly improves.
2 views 3 views 4 views 6 views
Refer to caption Refer to caption Refer to caption Refer to caption
12 views 24 views 36 views Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Additional synthetic-data ablation on the number of views used for training. With more input views, the visual quality improves, albeit hitting diminishing returns at 24 views.
Appendix F Ablation on FPS of Supporting RGB Frames
We ablate the number of supporting RGB frames used for training (Figs. 9 and 5). There is only a minimal difference in the results if we use only one RGB frame for reconstruction (0.5 FPS), compared to using 100 FPS RGB inputs. This indicates that our method does not depend on the RGB inputs much, using mostly the events. That could lead to a follow-up work that eliminates the RGB inputs.
Appendix G Ablations on View Count
When reducing the number of training views, we see that the model does not diverge even when using only three views (Figs. 10 and 6). However, we note that the accuracy of geometry increases significantly when using more views. We also used synthetic data to test more possible setups, up to 36 views. There is a clear improvement in PSNR with the increase in the number of training views as well. This confirms that our method indeed benefits from additional improvements to the setup, as stated in the conclusion.
Appendix H Ablation on Design Choices
We provide a detailed ablation study of our design choices in Tab. 4, listing quantitative results for all our synthetic scenes individually. Tab. 4 clearly shows that our full model performs best overall. We also provide a similar table for one of the real scenes in Tab. 3. SSIM and LPIPS also clearly favour the full model in that case.
FPS PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
0.5 $26.950$ $0.817$ $0.123$
1 $26.932$ $0.817$ $0.123$
2 $27.055$ $0.818$ $0.119$
5 $27.061$ $0.820$ $0.115$
10 $27.289$ $0.821$ $0.117$
20 $27.226$ $0.821$ $0.114$
30 $27.337$ $0.821$ $0.116$
50 $27.310$ $0.820$ $0.117$
100 $27.328$ $0.821$ $0.116$
Table 5: Quantitative study on the number of supporting RGB frames. By default, we use 5 FPS, the same frequency as our raw data. SSIM and LPIPS metrics favour values above or equal to 5 FPS, but do not drop too much with fewer frames, indicating that the model primarily uses event information and not RGB.
Input Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
2 $24.667$ $0.784$ $0.166$
3 $26.081$ $0.808$ $0.149$
4 $26.891$ $0.815$ $0.118$
5 $27.061$ $0.820$ $0.115$
Table 6: Quantitative study on the number of input views on the “Sword” real scene. All metrics clearly confirm the improvement with additional views.
2 v. 3 v. 4 v. 6 v. 12 v. 24 v. 36 v.
$16.05$ $28.97$ $27.90$ $30.88$ $32.19$ $33.36$ $33.27$
Table 7: PSNR scores on “Dress” with varying number of views. As stated in the conclusion, our method could easily benefit from an increase in the number of views or the resolution of the cameras, which is evident from the results. Bold indicates the two best-performing model results.
Appendix I Accumulation Stability
Figure 12: Demonstration of accumulation decay (Sec. 4.4). In a small patch of pixels (marked red on the right), we accumulate events for each pixel individually and show the resulting signals. Naive method (red) becomes unstable, as all pixels have completely different values at the end. In contrast, our proposed method with decay (blue) is stable: all pixels keep similar values at any time. Visually, the image with decay (right) has much fewer artefacts, too.
We have observed that accumulating long sequences of events leads to unstable results. This means that even in a pixel of constant brightness, spurious “noise” events will accumulate to ever-increasing deviations from the true brightness level. In this section, we prove the existence of this phenomenon analytically. We also show that our decay approach successfully limits this problem to a tolerable bound.
To see that noise events can destroy all information about the true brightness level, let us consider a pixel of constant brightness that, nevertheless, reports noise events with polarities $p_{i}\in\{-1,+1\}$. For the sake of simplicity, we assume that the $p_{i}$ are independent and identically distributed random variables with fixed, but arbitrary probabilities ${{q}^{+}}:=\mathbb{P}(p_{i}=+1)$ and ${{q}^{-}}:=\mathbb{P}(p_{i}=-1)$. The accumulated polarity after $n$ such events can then be expressed as the random variable
$E_{n}:=\sum\limits^{n}_{i=1}p_{i},$ (13)
which is a simplified version of Eq. 11. The expected value of $E_{n}$ is
$\mathbb{E}(E_{n})\stackrel{{\scriptstyle\text{Def.}\mathbb{E}}}{{=}}\sum\limits^{n}_{i=1}(+1)\cdot{{q}^{+}}+(-1)\cdot{{q}^{-}}=n\cdot({{q}^{+}}-{{q}^{-}}).$ (14)
If ${{q}^{+}}\neq{{q}^{-}}$, then $\lim\limits_{n\rightarrow\infty}\mathbb{E}(E_{n})=\pm\infty$, i.e. as one accumulates more and more noise, one can safely expect to drift arbitrarily far away from the true brightness level. But even when ${{q}^{+}}={{q}^{-}}$, which would make $\mathbb{E}(E_{n})=0$ for all $n$, the variance of (and therefore the confidence in) the accumulated polarity will decrease with growing $n$:
$\begin{array}[]{lcl}&\mathbb{V}(E_{n})\par\stackrel{{\scriptstyle\text{Def.}\mathbb{V}}}{{=}}&\mathbb{E}((E_{n}-\mathbb{E}(E_{n}))^{2})\\ &\stackrel{{\scriptstyle\text{Binom. thm.}}}{{=}}&\mathbb{E}(E_{n}^{2}-2E_{n}\cdot\mathbb{E}(E_{n})+\mathbb{E}(E_{n})^{2})\\ &\stackrel{{\scriptstyle\text{Additivity}\;\mathbb{E}}}{{=}}&\mathbb{E}(E_{n}^{2})-\mathbb{E}(E_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Def.}E_{n}}}{{=}}&\mathbb{E}\left(\sum\limits_{\begin{subarray}{c}1\leq i\leq n\\ 1\leq j\leq n\end{subarray}}p_{i}p_{j}\right)-\mathbb{E}(E_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Arithmetic}}}{{=}}&\mathbb{E}\left(\sum\limits_{i=1}p_{i}^{2}+2\cdot\sum\limits_{1\leq i<j\leq n}p_{i}p_{j}\right)-\mathbb{E}(E_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Additivity}\;\mathbb{E}}}{{=}}&\sum\limits_{i=1}\mathbb{E}(p_{i}^{2})+2\cdot\sum\limits_{1\leq i<j\leq n}\mathbb{E}(p_{i}p_{j})-\mathbb{E}(E_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Def.}\mathbb{E}}}{{=}}&n\cdot({{q}^{+}}+{{q}^{-}})\\ &&+2\cdot\sum\limits_{1\leq i<j\leq n}{{q}^{+}}^{2}-2{{q}^{+}}{{q}^{-}}+{{q}^{-}}^{2}\\ &&-n\cdot({{q}^{+}}-{{q}^{-}})\\ &\stackrel{{\scriptstyle\begin{subarray}{c}\text{Binom. thm.}\\ \text{Arithmetic}\end{subarray}}}{{=}}&2({{q}^{+}}-{{q}^{-}})^{2}\cdot\sum\limits_{i=1}^{n-1}i+2n{{q}^{-}}\\ &\stackrel{{\scriptstyle\begin{subarray}{c}\text{Gauß sum}\\ \text{Arithmetic}\end{subarray}}}{{=}}&n\cdot(n-1)\cdot({{q}^{+}}-{{q}^{-}})^{2}+2n{{q}^{-}}.\par\end{array}$ (15)
This term can exceed any arbitrary bound if one lets $n$ grow large enough: If ${{q}^{-}}=0$, then ${{q}^{+}}=1$ and $\mathbb{V}(E_{n})=n^{2}-n$. If ${{q}^{+}}-{{q}^{-}}=0$, then ${{q}^{+}}={{q}^{-}}=\frac{1}{2}$ and $\mathbb{V}(E_{n})=n$. In all other cases we have ${{q}^{-}}>0$ and $({{q}^{+}}-{{q}^{-}})^{2}>0$ and thus also $\lim\limits_{n\rightarrow\infty}\mathbb{V}(E_{n})=+\infty$.
This shows that the noise events will drown out all information about the actual brightness level, if one just waits long enough. Since this effect would deteriorate not only background pixels (which should not cause any events), but also foreground pixels (where noise is interleaved with legitimate events), we need to achieve $\lim\limits_{n\rightarrow\infty}\mathbb{V}(E_{n})\in\mathbb{R}$ and thus introduce decay: When accumulating polarities, we value older events less than younger events:
$\hat{E}_{n}:=\sum\limits^{n}_{i=1}p_{i}\cdot b^{n-i},$ (16)
where the decay strength $b:=0.93$ was empirically found to be a useful value. Now we have:
$\begin{array}[]{lcl}\mathbb{E}(\hat{E}_{n})\par&\stackrel{{\scriptstyle\text{Additivity}\;\mathbb{E}}}{{=}}&\sum\limits_{i=1}^{n}\mathbb{E}(p_{i})\cdot b^{n-i}\\ &\stackrel{{\scriptstyle\text{Def.}\;\mathbb{E}}}{{=}}&({{q}^{+}}-{{q}^{-}})\cdot\sum\limits_{i=1}^{n}b^{n-i}\\ &\stackrel{{\scriptstyle\text{Arithmetic}}}{{=}}&({{q}^{+}}-{{q}^{-}})\cdot\sum\limits_{i=0}^{n-1}b^{i}\\ &\stackrel{{\scriptstyle\text{Geometric sum}}}{{=}}&({{q}^{+}}-{{q}^{-}})\cdot\frac{b^{n}-1}{b-1}.\par\end{array}$ (17)
Since $\lim\limits_{n\rightarrow\infty}b^{n}=0$, we have that $\lim\limits_{n\rightarrow\infty}\mathbb{E}(\hat{E}_{n})$ is finite. In a derivation similar to Eq. 15, we obtain
$\begin{array}[]{lcl}\mathbb{V}(\hat{E}_{n})\par&\stackrel{{\scriptstyle\begin{subarray}{c}\text{Def.}\mathbb{V}\\ \text{Binom. thm.}\\ \text{Additivity}\mathbb{E}\end{subarray}}}{{=}}&\mathbb{E}(\hat{E}_{n}^{2})-\mathbb{E}(\hat{E}_{n})^{2}\\ &\stackrel{{\scriptstyle\begin{subarray}{c}\text{Def.}\;\hat{E}_{n}\\ \text{Arithmetic}\\ \text{Additivity}\mathbb{E}\end{subarray}}}{{=}}&\sum\limits_{i=1}^{n}\mathbb{E}(p_{i}^{2})\cdot b^{2(n-i)}\\ &&+2\cdot\sum\limits_{1\leq i<j\leq n}\mathbb{E}(p_{i}p_{j})\cdot b^{n-i}\cdot b^{n-j}\\ &&-\mathbb{E}(\hat{E}_{n})^{2}\\ &\stackrel{{\scriptstyle\begin{subarray}{c}\text{Def.}\mathbb{E}\\ \text{Binom. thm.}\\ \text{Arithmetic}\end{subarray}}}{{=}}&({{q}^{+}}+{{q}^{-}})\cdot\sum\limits_{i=0}^{n-1}(b^{2})^{i}\\ &&+2\cdot({{q}^{+}}-{{q}^{-}})^{2}\cdot\sum\limits_{i=0}^{n-1}b^{i}\cdot\sum\limits_{j=0}^{i-1}b^{j}\\ &&-\mathbb{E}(\hat{E}_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Geometric sum}}}{{=}}&({{q}^{+}}+{{q}^{-}})\cdot\frac{b^{2n}-1}{b^{2}-1}\\ &&+2\cdot({{q}^{+}}-{{q}^{-}})^{2}\cdot\sum\limits_{i=0}^{n-1}b^{i}\cdot\frac{b^{i}-1}{b-1}\\ &&-\mathbb{E}(\hat{E}_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Arithmetic}}}{{=}}&({{q}^{+}}+{{q}^{-}})\cdot\frac{b^{2n}-1}{b^{2}-1}\\ &&+2\cdot\frac{({{q}^{+}}-{{q}^{-}})^{2}}{b-1}\cdot\left(\sum\limits_{i=0}^{n-1}(b^{2})^{i}-\sum\limits_{i=0}^{n-1}b^{i}\right)\\ &&-\mathbb{E}(\hat{E}_{n})^{2}\\ &\stackrel{{\scriptstyle\text{Geometric sum}}}{{=}}&({{q}^{+}}+{{q}^{-}})\cdot\frac{b^{2n}-1}{b^{2}-1}\\ &&+2\cdot\frac{({{q}^{+}}-{{q}^{-}})^{2}}{b-1}\cdot(\frac{b^{2n}-1}{b^{2}-1}-\frac{b^{n}-1}{b-1})\\ &&-\mathbb{E}(\hat{E}_{n})^{2}.\par\end{array}$ (18)
All summands in the last term of Eq. 18 converge to a finite number as $n$ grows larger, so $\lim\limits_{n\rightarrow\infty}\mathbb{V}(\hat{E}_{n})\in\mathbb{R}$. This shows that decay is effectively bounding the deviation from the true brightness level to a finite error, even for arbitrary numbers of noise events.
Fig. 12 (right) shows that after applying decay, accumulated events become clear, even though they were unrecognisable without decay, proving the effectiveness of our strategy. Additionally, we take a small patch of that view and plot the accumulated values with and without decay in Fig. 12 (left): We see that, as predicted analytically, without decay, the accumulation grows beyond all bounds due to the noise, while with decay, it remains in a constant range, exactly as the signal should be.
The smaller one chooses $b$, the smaller $\lim\limits_{n\rightarrow\infty}\mathbb{V}(\hat{E}_{n})$ becomes and the faster it converges. However, since decay affects not only noise events but also legitimate ones, we have to explain why it does not distort the actual signal too much. To see this, consider a single pixel: As long as the foreground is not moving through this pixel, it shows a constant level of background brightness. All events it emits are noise, and can safely be suppressed by decay (because the accumulated polarity is supposed to be zero). When the foreground enters the pixel, it will trigger a number of legitimate events. As long as these are recent, the decay will not suppress them too much, so $\hat{E}$ is approximately at the foreground brightness level. Assuming that $b$ is chosen suitably, motion is usually fast enough that the object will leave the pixel again before decay “undoes” the entry events completely. Leaving the pixel again triggers a set of legitimate events that are first sufficiently recent not to be suppressed by damping. As they move further into the past, so do the entry events, so $\hat{E}$ will then correctly approximate the background brightness level again. Of course, decay does still negatively impact the legitimate events (especially when motion is occasionally slower than what $b$ was tuned for, such that the foreground is inside the pixel for longer durations), but the underlying MLP, supervised by all our losses, can compensate for that sufficiently. The ablations in Tabs. 3 and 4 show that SSIM scores, LPIPS and visual results are indeed improved by our decay technique.